# Spacetime games subsume causal contextuality scenarios

### v3

@@ 4.4 Causal contextuality scenarios
In this paper, we consider “clean” scenarios in the sense that there are no unused measurement settings, i.e., settings that cannot be enabled under any circumstances under the considered cover. If a causal contextuality scenario contains an unused setting, then it can be removed without affecting its semantics. This is analogous to our assumption that spacetime games do not have unused information sets.
Furthermore, this paper restricts its attention to covers that satisfy the antichain property. It is, however, possible to extend the mapping naturally even if the antichain property does not hold (e.g., by allowing unused actions at Bob’s nodes).
4.5 Natural cover
@@
Each spacetime game (not necessarily alternating) has a natural cover that corresponds to all combinations of measurements that are performed in one of the complete histories:
@@
Given a spacetime game $\mathcal{G}$ with nature player Alfred, the natural cover $\mathcal{C}$ is defined as follows: For each outcome $z\in\mathcal{Z}$, we take the subset $C_{z}=\{i\in\textsf{domain}(z)|\rho(i)=\text{Alfred}\}$ of its domain of definition with only and exactly the information sets that are played by Alfred. Then the natural cover is $\mathcal{C}=\{C_{z}|z\in\mathcal{Z}\}$. Note that duplicates are implicitly eliminated because of the definition of a set.
@@
The natural cover of the spacetime game shown in Figure 5 and Figure 6 is $\{\{X\},\{YW\},\{YZ\}\}$.
This natural cover will be used to define the corresponding measurement scenario for any alternating spacetime game.
4.6 The mapping
4.6.1 Causal contextuality scenario associated with an alternating spacetime game
Any alternating spacetime game can be associated with a causal contextuality scenario as follows. As explained in the previous tabular correspondence, the general idea of the overall mapping is that:
  • Alfred’s information sets are measurements;
  • Bob’s (singleton) information sets are the left-hand sides of the enabling relations of the causal contextuality scenario;
@@ 4.6.1 Causal contextuality scenario associated with an alternating spacetime game
Definition 7 (Causal contextuality scenario associated with an alternating spacetime game).
Given an alternating spacetime game $\mathcal{G}$ with nature player $\text{Alfred}\in\mathcal{P}$, we define the equivalent causal contextuality scenario $\Sigma(\mathcal{G})=(X,O,\vdash,\mathcal{C})$ as follows:
  • $X=\{i\in\mathcal{I}|\rho(i)=\text{Alfred}\}$
@@
  • for each $x\in X$, and for each $P\in\textsf{predecessors}(x)$, we define that $t_{P}\vdash x$ where
    $t_{P}=\{(\iota(M),\sigma(M,P))|M\smile P\}$
@@
For it to be a valid causal contextuality scenario, we must prove that the enabling relations have consistent left-hand sides.
@@
The causal contextuality scenario associated with an alternating spacetime game always has consistent sets of events on the left-hand sides of enabling relations.
@@
AB1 implies that two nodes in $y$, which is played by Alfred, have the same labels and destination nodes. As a consequence, if we have two nodes $M,N\in y$, it follows that $\sigma(M,P)=\sigma(N,P)$, leading to only one assignment for $t(y)$. ∎
@@
It follows from the lemma, given some $x\in X$ and one of its predecessors $P$, that $\sigma(M,P)=t_{P}(y)$ for any $M\in y$. See also Figure 9 for a visual of this pattern.
Figure 9: The graph pattern implied by the definition of $t\vdash x$ in the mapping from games to causal contextuality scenarios. We show the pattern for two events in $t=\{(y,t(y)),(z,t(z))\}$ two contexts for $y$ and $z$, and four contexts (among which two under causal bridge $t$) for $x$, but this extends to any number of events and any domain size.
@@
$P\mapsto t_{P}$ is a bijection between the set of nodes played by Bob and the left-hand sides of the mapped enabling relations. More precisely, for any $x\in X$, the left-hand sides of the enabling relations to $x$ in the causal contextuality scenario are in bijection with the predecessors of $x$ in the alternating spacetime game.
@@
Every $P$ leads to at least one enabling relation with left-hand-side $t_{P}$ because it has at least one child (property EVEN). It follows from the property AB2 that $t$ is injective. Since the codomain is defined as the range, the mapping is also surjective. ∎
We use the abbreviation $\textsf{enabled}(t)=\{x\in X|t\vdash x\}$ and use this notation in particular for restricting the cover $\mathcal{C}$ to the “local cover at t” denoted $\mathcal{C}_{t}$ which we define as
Note that $\mathcal{C}_{t}$ cannot be empty because there are no unused settings, and it does not contain the empty set.
4.6.2 Alternating spacetime game associated with a causal contextuality scenario

### v4

@@ 3.2.1 Definition
In this paper, we consider “clean” scenarios in the sense that there are no unused measurement settings, i.e., settings that cannot be enabled under any circumstances under the considered cover. This amounts to saying that $\bar{\tau}(x)$ is always a consistent set of events. If a causal contextuality scenario contains an unused setting, then it can be removed without affecting its semantics. This is analogous to our assumption that spacetime games do not have unused information sets.
Definition 10 (Local cover restriction).
We define the local cover restriction $\mathcal{C}_{t}$ at an enabling set of events $t$ the set of the facets of the restriction to $\textsf{enabled}(t)$ of the simplicial map associated with $\mathcal{C}$.
Remark 5.
We have
$\forall x\in X,\exists C\in\mathcal{C}_{\tau(x)},x\in C$
[11] mentions that one can define “good covers” such that the strategy presheaf fulfills the sheaf property, but a definition thereof is not known to us as we could not find it in publicly available repositories. Since this nice property also characterizes those covers that make a causal contextuality scenario equivalent in structure to an alternating spacetime game, we give a definition of such covers that leads to the categorical equivalence of the two structures. We use the terminology “causally-secured cover” which was used in recent presentations by [11], with the warning that this definition may differ from future publications from the causal contextuality scenario research team.
Definition 11 (Causally-secured cover).
A causal contextuality scenario $\Gamma=(X,O,\vdash,\mathcal{C})$ is said to have a causally secured cover if
  • Facet membership propagates to all enabling measurements:
    $\forall C\in\mathcal{C},\forall C\,\text{support}(\tau(x))\subseteq C$
  • Two measurements with inconsistent enabling conditions never belong to the same facet:
    $\forall x,y\textsf{ such that }(\bar{\tau}(x)\cup\bar{\tau}(y))\text{ is inconsistent },\forall C\in\mathcal{C},\{x,y\}\not\subseteq C$
  • $\mathcal{C}$ is the maximum among all covers $\mathcal{C}^{\prime}$ that have the same local cover restrictions ($\forall t,\mathcal{C}_{t}^{\prime}=\mathcal{C}_{t}$) and that fulfill the other two criteria [^15] This criterion is akin to some form of non-signaling in the available choices of local covers at any spacetime location..
Remark 6.
By definition, it means that a causally-secured cover is fully specified by all its local cover restrictions.
Remark 7.
It follows from the maximality of a causally secured cover that the local cover restrictions can be characterized with
that is, it is not necessary to explicitly eliminate non-maximal faces.
3.2.2 Category of causal contextuality scenarios
We now extend the morphism of [23], which was defined on flat scenarios, to acyclic causal contextuality scenarios with causally secured covers and unique causal bridges (thereafter scenario), i.e., to make it compatible with the enabling relation. We are not aware of a follow-up to [23] that defines the morphisms in the causal case.
Definition 12 (Scenario morphism).
Consider two scenarios $\Gamma=(X,O,\vdash,\mathcal{C})$ and $\Gamma^{\prime}=(X^{\prime},O^{\prime},\vdash^{\prime},\mathcal{C}^{\prime})$. A morphism $\gamma:\Gamma^{\prime}\rightarrow\Gamma$ is a tuple $(\pi^{\prime},\alpha=\{\alpha_{x}\}_{x\in X})$ where:
  • $\pi^{\prime}$ is a function from $\mathcal{X}$ to $\mathcal{X^{\prime}}$ that entails a simplicial map from $\mathcal{C}$ to $\mathcal{C^{\prime}}$.
  • for each $x\in X$, $\alpha_{x}$ is a function from $O^{\prime}_{\pi(x)}$ to $O_{x}$.
With the following structural constraints:
  1. 1.
    $\pi^{\prime}$ preserves causal bridges:
    $\forall x,y,\tau(\pi^{\prime}(x))=\tau(\pi^{\prime}(y))\implies\tau(x)=\tau(y)$
  2. 2.
    Measurement enablement is not precluded by $\pi^{\prime}$, i.e., for any $x\in X$, any $y^{\prime}$ in the support of $\tau(\pi^{\prime}(x))$ must be in the image of $\pi^{\prime}$.
    $\forall x\in X,\forall y^{\prime}\in\textsf{support}(\tau(\pi^{\prime}(x)),\exists y\in X,\pi^{\prime}(y)=y^{\prime}$
  3. 3.
    For any $x\in X$ and any $y\in X$ such that $\pi^{\prime}(y)\in\textsf{support}(\tau(\pi^{\prime}(x))$:
    $\alpha_{y}\left[\tau(\pi^{\prime}(x))(\pi^{\prime}(y)\right]=\tau(x)(y)$
Remark 8.
Rules 2 and 3 imply that, for any $x\in X$, if $\tau(x)=\{(x_{i},o_{i})\}_{i}$, then $\tau(\pi^{\prime}(x))$ can be written in the form $\{(\pi^{\prime}(x_{i}),o_{i}^{\prime})\}_{i}$ where:
$\forall i,\alpha_{x_{i}}(o_{i}^{\prime})=o_{i}$
Because all enabling relations must have a consistent set of events as their left-hand-side, the existence of an enabling relation of that form also implies in rule 1:
$\forall i,j,\pi^{\prime}(x_{i})=\pi^{\prime}(x_{j})\implies o_{i}^{\prime}=o_{j}^{\prime}$
Remark 9.
Rule 3 also implies that if $t\vdash x$ and $t\vdash y$ then $\tau(\pi^{\prime}(x))=\tau(\pi^{\prime}(y))$ because the form of $\tau(\pi^{\prime}(x))$ depends only on $\tau(x)$ and otherwise not on the particular choice of $x$. Then there is $t^{\prime}$ such that $t^{\prime}\vdash\pi^{\prime}(x)$ and $t^{\prime}\vdash\pi^{\prime}(y)$.
This implies that the causal bridge conservation implication is in fact an equivalence:
$\forall x,y,\tau(\pi^{\prime}(x))=\tau(\pi^{\prime}(y))\iff\tau(x)=\tau(y)$
There is a scenario morphism that turns any scenario into itself:
Definition 13 (Identity morphism for a scenario with a causally secured cover).
Consider a scenario $\Gamma=(X,O,\vdash,\mathcal{C})$ The identity morphism $\text{Id}_{\Gamma}:\Gamma\rightarrow\Gamma$ is the tuple $(\pi^{\prime},\alpha)$ where:
  • $\pi^{\prime}=\text{Id}_{X}$.
  • $\alpha_{x}=\text{Id}_{O_{x}}$ for each $x\in X$.
Lemma 6.
The identity morphism for any scenario is a morphism.
Proof.
It is straightforward to verify that the identity morphism follows the required structural constraints as they transform into tautologies. ∎
Lemma 7 (Left and right unit laws).
Consider two scenarios with no cycles, unique causal bridges, and a causally secured cover $\Gamma=(X,O,\vdash,\mathcal{C})$ and $\Gamma^{\prime}=(X^{\prime},O^{\prime},\vdash^{\prime},\mathcal{C}^{\prime})$ as well as a morphism $\gamma:\Gamma^{\prime}\rightarrow\Gamma$ with components $(\pi^{\prime},\alpha)$. Then $\gamma=\text{Id}_{\Gamma}\circ\gamma=\gamma\circ\text{Id}_{\Gamma^{\prime}}$
Proof.
It follows from applying the left and right unit laws on the individual components. ∎
And morphisms can be composed like so:
Definition 14 (Composition of morphisms on scenarios).
Consider three scenarios $\Gamma$, $\Gamma^{\prime}$, and $\Gamma^{\prime\prime}$ and two morphisms $\mu_{1}:\Gamma^{\prime}\rightarrow\Gamma$ with components $(\pi_{1}^{\prime},\alpha_{1})$ and $\mu_{2}:\Gamma^{\prime\prime}\rightarrow\Gamma^{\prime}$ with components $(\pi_{2}^{\prime\prime},\alpha_{2}^{\prime})$. The morphism $\mu_{3}=\mu_{1}\circ\mu_{2}:\Gamma^{\prime\prime}\rightarrow\Gamma$ is defined as the tuple $(\pi_{3}^{\prime\prime},\alpha_{3})$ where:
  • $\pi_{3}^{\prime\prime}=\pi_{2}^{\prime\prime}\circ\pi_{1}^{\prime}$.
  • $\alpha_{3,x}=\alpha_{1,x}\circ\alpha_{2,\pi_{1}^{\prime}(x)}^{\prime}$ for every $x\in X$.
Lemma 8.
The result of morphism decomposition as defined above is a morphism.
Proof.
Let $x,y\in X$ such that
$\tau(\pi_{3}^{\prime\prime}(x))=\tau(\pi_{3}^{\prime\prime}(y))$
Then we have
$\tau(\pi_{2}^{\prime\prime}(\pi_{1}^{\prime}(x)))=\tau(\pi_{2}^{\prime\prime}(\pi_{1}^{\prime}(y)))$
Since $\mu_{2}$ is a morphism:
$\tau(\pi_{1}^{\prime}(x))=\tau(\pi_{1}^{\prime}(y))$
Since $\mu_{1}$ is a morphism:
$\tau(x)=\tau(y)$
which proves the first rule.
Let $x\in X$ and $y^{\prime\prime}\in\textsf{support}(\tau(\pi_{3}^{\prime\prime}(x))$.
It follows that $y^{\prime\prime}$ is also in the support of $\tau(\pi_{2}^{\prime\prime}(\pi_{1}^{\prime}(x))$).
Because $\mu_{2}$ is a morphism, there is some $y^{\prime}\in X^{\prime}$ such that $\pi_{2}^{\prime\prime}(y^{\prime})=y^{\prime\prime}$.
Rule 3 applied $\mu_{2}$, $\pi_{1}^{\prime}(x)$ and $y^{\prime}$ implies that $y^{\prime}$ is in the support of $\tau(\pi_{1}^{\prime}(x))$.
Since $\mu_{1}$ is a morphism, there is some $y\in X$ such that $\pi_{1}^{\prime}(y)=y^{\prime}$.
It follows that $\pi_{3}^{\prime\prime}(y)=\pi_{2}^{\prime\prime}(\pi_{1}^{\prime}(y))=\pi_{2}^{\prime\prime}(y^{\prime})=y^{\prime\prime}$.
This proves the second rule.
Let there be $x\in X$ and $y\in X$ so that $\pi_{3}^{\prime\prime}(y)\in\textsf{support}(\tau(\pi_{3}^{\prime\prime}(x))$.
Denote $y^{\prime}=\pi_{1}^{\prime}(y)$ and $x^{\prime}=\pi_{1}^{\prime}(x)$.
Then $\pi_{2}^{\prime\prime}(y^{\prime})=\pi_{3}^{\prime\prime}(y)\in\textsf{support}(\tau(\pi_{3}^{\prime\prime}(x))=\textsf{support}(\tau(\pi_{2}^{\prime\prime}(x^{\prime}))$.
Because $\mu_{2}$ is a morphism, we have
$\alpha_{2,y^{\prime}}(\tau(\pi_{2}^{\prime\prime}(x^{\prime}))(\pi_{2}^{\prime\prime}(y^{\prime}))=\tau(x^{\prime})(y^{\prime})=\tau(\pi_{1}^{\prime}(x))(\pi_{1}^{\prime}(y))$
Then $\pi_{1}^{\prime}(y)$ is in the support of $\tau(\pi_{1}^{\prime}(x))$ and because $\mu_{1}$ is a morphism, we have
$\alpha_{1,y}(\tau(\pi_{1}^{\prime}(x))(\pi_{1}^{\prime}(y))=\tau(x)(y)$
Combining the two, we get:
$\alpha_{1,y}(\alpha_{2,\pi_{1}^{\prime}(y)}(\tau(\pi_{2}^{\prime\prime}(\pi_{1}^{\prime}(x)))(\pi_{2}^{\prime\prime}(\pi_{1}^{\prime}(y))))=\tau(x)(y)$
Which is to say:
$\alpha_{3,y}(\tau(\pi_{3}^{\prime\prime}(x))(\pi_{3}^{\prime\prime}(y)))=\tau(x)(y)$
which proves the third rule.
Thus, $\gamma_{3}$ is a morphism. ∎
Lemma 9 (Associativity of morphism composition).
$\circ$ as defined on scenario morphisms is associative.
Proof.
If we take four games $\Gamma$, $\Gamma^{\prime}$, $\Gamma^{\prime\prime}$, and $\Gamma^{\prime\prime\prime}$ and three morphisms $\gamma_{1}:\Gamma^{\prime}\rightarrow\Gamma$ with components $(\pi_{1}^{\prime},\alpha_{1})$, $\gamma_{2}:\Gamma^{\prime\prime}\rightarrow\Gamma^{\prime}$ with components $(\pi_{2}^{\prime\prime},\alpha_{2}^{\prime})$, and $\gamma_{3}:\Gamma^{\prime\prime\prime}\rightarrow\Gamma^{\prime\prime}$ with components $(\pi_{3}^{\prime\prime\prime},\alpha_{3}^{\prime\prime})$ then $(\gamma_{3}\circ\gamma_{2})\circ\gamma_{1}$ is by definition:
  • $\pi_{4}^{\prime\prime}=(\pi_{3}^{\prime\prime\prime}\circ\pi_{2}^{\prime\prime})\circ\pi_{1}^{\prime}$.
  • $\alpha_{4,x}=\alpha_{1,x}\circ(\alpha_{2,\pi_{1}^{\prime}(x)}^{\prime}\circ\alpha_{3,\pi_{2}^{\prime\prime}(\pi_{1}^{\prime}(x))}^{\prime\prime})$.
which is the same as $\gamma_{3}\circ(\gamma_{2}\circ\gamma_{1})$:
  • $\pi_{4}^{\prime\prime}=\pi_{3}^{\prime\prime\prime}\circ(\pi_{2}^{\prime\prime}\circ\pi_{1}^{\prime})$.
  • $\alpha_{4,x}=(\alpha_{1,x}\circ\alpha_{2,\pi_{1}^{\prime}(x)})^{\prime}\circ\alpha_{3,\pi_{2}^{\prime\prime}(\pi_{1}^{\prime}(x))}^{\prime\prime}$.
Remark 10.
A morphism $\gamma:\Gamma^{\prime}\rightarrow\Gamma$ with components $(\pi^{\prime},\alpha)$ is a section if $\pi^{\prime}$ is a section and $\alpha$ is a retraction. It is a retraction if $\pi^{\prime}$ is a retraction and $\alpha$ is a section.
Theorem 10 (Category of scenarios).
The class of all causal contextuality scenarios with no cycles, unique causal bridges, and a causally-secured cover, endowed with the morphisms as defined above, is a category. We denote it Scenario.
Proof.
This follows from the left and right unit laws and the associativity of morphism composition. ∎
3.3 Categorical equivalence
Our goal is to prove the following theorem:
Theorem 11.
The categories Game and Scenario are equivalent:
[…]
For this, we need to define a functor that is essential surjective and fully faithful.
3.3.1 Game-Scenario functor
We now define the functor $F:\textbf{Game}\rightarrow\textbf{Scenario}$.
First, we map the objects. Any alternating spacetime game can be associated with a scenario as follows. As explained in the previous tabular correspondence, the general idea of the overall mapping is that:
  • Alfred’s information sets are measurements;
  • Bob’s (singleton) information sets are the left-hand sides of the enabling relations of the scenario;
@@ 3.3.1 Game-Scenario functor
In order to define the cover of the associated scenario, we first show that each spacetime game (not necessarily alternating) has a natural cover that corresponds to the facets of all combinations of measurements that are performed in one of the complete histories:
@@
Given a spacetime game $\mathcal{G}$ with nature player Alfred, the natural cover $\mathcal{C}$ is defined as follows: For each outcome $z\in\mathcal{Z}$, we take the subset $C_{z}=\mathcal{I}_{A}\cap\textsf{support}(z)$ of its domain of definition with only and exactly the information sets that are played by Alfred. Then the natural cover is $\mathcal{C}=\textsf{facets}(\{C_{z}|z\in\mathcal{Z}\}$) where facets eliminates the non-maximal elements, i.e., returns the facets of the simplicial complex obtained via the downward closure of its argument. Note that duplicates are implicitly eliminated because of the definition of a set.
The natural cover associated with a game thus corresponds to the facets of a simplicial complex whose vertices are the universe information sets.
@@
The natural cover of the spacetime game shown in Figure 5 and Figure 6 is $\{\{X\},\{YW\},\{YZ\}\}$ (with $\{Y\}$ eliminated by facets).
This natural cover will now be used to define the corresponding measurement scenario for any alternating spacetime game.
@@
Definition 16 (Scenario associated with an alternating spacetime game).
Given an alternating spacetime game $\mathcal{G}$ which is an object in the Game category, we define the equivalent scenario $F(\mathcal{G})=(X,O,\vdash,\mathcal{C})$ as follows:
  • $X=\mathcal{I}_{A}$
@@
  • for each $x\in X$, with $t=\textsf{parent}(x)$, we define that $\hat{t}\vdash x$ where
    $\hat{t}=\{(\iota(m),\sigma(m,t))|m\in\mathcal{N}_{A},m\smile t\}$
@@
For it to be a valid scenario, we must prove that the enabling relations have consistent left-hand sides.
@@
The scenario associated with an alternating spacetime game always has consistent sets of events on the left-hand sides of enabling relations.
@@
AB1 implies that two nodes in $y$, which is played by Alfred, have the same labels and destination nodes. As a consequence, if we have two nodes $m,n$ such that $\iota(m)=\iota(n)=y,m\smile x,n\smile x$, i.e., $m,n\in y$, it follows that $\sigma(m,t)=\sigma(n,t)$, leading to only one assignment for $t(y)=\sigma(y,t)$. ∎
@@
It follows from the lemma, given some $x\in X$ and its parent $t$, that $\sigma(m,t)=\hat{t}=\sigma(y,t)$ for any $m\in y$. This means in particular that we can write $\hat{t}$ in this equivalent form:
$\hat{t}=\{(y,\sigma(y,t))|y\in\mathcal{I}_{A},y\smile t\}$
See also Figure 11 for a visual of this pattern.
We must also prove that the functor maps games with scenarios that have a causally secured cover.
@@
If $\mathcal{G}\in ob(\textbf{Game})$, then $F(\mathcal{G})$ has a causally secured cover.
@@
Let us start with the first criterion. Let $F(\mathcal{G})=(X,O,\vdash,\mathcal{C})$. Let $C\in\mathcal{C}$. Since $\mathcal{C}$ is a natural cover of $\mathcal{G}$, it follows that there exists a complete history $z$ such that $C_{z}=C$. By definition, it follows that $C\subseteq\{x\in\textsf{support}(z)|\rho(x)=\text{Alfred}\}=\textsf{support}(z)\cap\mathcal{I}_{A}$. Let $x\in C$. It follows that $x\in\textsf{support}(z)$ and that $\rho(x)=\text{Alfred}$. By BA1, it follows that $x$ has exactly one parent node t and thus there is, by definition of $F(\mathcal{G})$, we have
$\hat{t}=\{(y,\sigma(y,t))|y\smile t\}$
and $\hat{t}\vdash x$. Furthermore, $t$ must be in support of the same history $z$ by definition of the history. Let $y$ be in the support of $\hat{t}$. Thus, we have $z(y)=\sigma(y,t)=\hat{t}(y)$. By definition of the history, $y$ must also be in the support of the history. Since $\rho(y)=\text{Alfred}$, it follows that $y\in C_{z}$. This is true for any $y$, so that $\textsf{support}(\tau(x))\subseteq C_{z}=C$. This is true for any C, so that $\mathcal{C}$ is a causally secured cover and $F(\mathcal{G})\in\text{obj}(\textbf{Scenario)}$).
For the second criterion: if two measurements have incompatible enabling relations, then they cannot both appear in the support the same history because this would cause an information set in the past to be associated with two distinct values, which is not possible by definition of a history.
For the third criterion: assume the natural cover $\mathcal{C}$ is not a maximum. Then there is another cover $\mathcal{D}$ that has the same local cover restrictions and that has a facet not contained in any facet of $\mathcal{C}$. Let us order the information sets in this facet as a linearization of the game (that is, in a linear order compatible with the DAG of the game) and consider the first information set $x$ in this facet such that the set of previous information sets is still included in a facet $C$ of $\mathcal{C}$ that is the support of some history $z$. This information set has a parent information set $t$ played by Bob corresponding to some enabling relations. But $x$ is a vertex in the local cover restriction of $\mathcal{C}$ at $t$ (which is the same as the local cover restriction of $\mathcal{D}$ at $t$ by assumption). This implies that $z$ is an incomplete history as it contains the support of $t$ but not $x$. This contradicts the assumption and thus, $\mathcal{C}$ is a maximum among all causally-secured covers with the same local cover restrictions. ∎
Figure 11: The graph pattern implied by the definition of $t\vdash x$ in the mapping from games to scenarios. We show the pattern for two events in $t=\{(y,t(y)),(z,t(z))\}$ two contexts for $y$ and $z$, and four contexts (among which two under causal bridge $t$) for $x$, but this extends to any number of events and any domain size.
Lemma 14.
$t\mapsto\hat{t}$ is a bijection between the set of nodes played by Bob and the left-hand sides of the mapped enabling relations. More precisely, for any $x\in X$, the left-hand sides of the enabling relations to $x$ in the scenario are in bijection with the predecessors of $x$ in the alternating spacetime game. When it is clear from the context, we will write $t=\hat{t}$.
Proof.
@@
Every $t$ leads to at least one enabling relation with left-hand-side $\hat{t}$ because it has at least one child (property EVEN). It follows from property AB2 that $t\mapsto\hat{t}$ is injective. Since the codomain is defined as the range, the mapping is also surjective. ∎
We now move on to how the functor maps the morphisms.
Definition 17.
Let $\mathcal{G},\mathcal{G^{\prime}}\in\textsf{ob}(\textbf{Game})$ (with their components having the usual notations) and let $\gamma:\mathcal{G^{\prime}}\rightarrow{G}$ be a morphism with components $(\nu^{\prime},\beta)$. Denoting the components $F(\mathcal{G})$ as $(X,O,\vdash,\mathcal{C})$ and those of $F(\mathcal{G^{\prime}})$ as $(X^{\prime},O^{\prime},\vdash^{\prime},\mathcal{C}^{\prime})$, we define the morphism $F(\gamma):F(\mathcal{G^{\prime}})\rightarrow F(\mathcal{G})$ as follows.
  • $\pi^{\prime}=\nu^{\prime}$ seeing $\nu^{\prime}$ as acting on information sets. This is mathematically consistent because $X=\mathcal{I}_{A}$ and $X^{\prime}=\mathcal{I}_{A}^{\prime}$, so that $\pi^{\prime}$ and $\nu^{\prime}$ have the same domain and codomain.
  • $\alpha=\beta$. This is also consistent because $\beta$ is a family of maps indexed by $X$ which is the same as $\mathcal{I}_{A}$, and $\beta_{x}$ for $x\in\mathcal{I}_{A}$ is a map from $\chi(\pi^{\prime}(x))\rightarrow\chi(x)$ which is the same as from $O_{\pi^{\prime}(x)}^{\prime}\rightarrow O_{x}$ by definition of F.
Lemma 15.
$F(\gamma)$ as defined is a scenario morphism.
Proof.
We need to check that each one of the three constraints applies.
Let there be $x,y\in X$ such that $\tau(\pi^{\prime}(x))=\tau(\pi^{\prime}(y))$.
By Lemma 14, $\tau(\pi^{\prime}(x))=\tau(\pi^{\prime}(y))$ correspond to one same node $t^{\prime}\in\mathcal{N}_{B}$ and by definition of $F$, this node is the parent of both $\pi^{\prime}(x)$ and of $\pi^{\prime}(y)$.
Thus
$\textsf{parent}(\pi^{\prime}(x))=\textsf{parent}(\pi^{\prime}(y))$
By definition of F, $\pi^{\prime}(x)=\nu^{\prime}(x)$ and $\pi^{\prime}(y)=\nu^{\prime}(y)$. Thus,
$\textsf{parent}(\nu^{\prime}(x))=\textsf{parent}(\nu^{\prime}(y))$
Because $\gamma$ is a morphism it follows that
$\textsf{parent}(x)=\textsf{parent}(y)$
It follows that $\tau(x)=\tau(y)$ by definition of $F$.
This proves that the first constraint holds.
Let there be $x\in X$ and $y^{\prime}\in\textsf{support}(\tau(\pi^{\prime}(x)))$.
It follows that $y^{\prime}\smile\textsf{parent}(\pi^{\prime}(x))$ by definition of $F$ and thus $y^{\prime}\smile\textsf{parent}(\nu^{\prime}(x))$ also by definition of $F$.
$\textsf{parent}(\nu^{\prime}(x))\in\mathcal{D}_{\nu}$ by definition of $\mathcal{D}_{\nu}$. Because $\gamma$ is a morphism,
$\exists y\in\mathcal{I}_{A},\nu^{\prime}(y)=y^{\prime}$
from which it follows by definition of $F$:
$\exists y\in\mathcal{I}_{A},\pi^{\prime}(y)=y^{\prime}$
This proves the second constraint.
Let there be $x,y\in X$ such that $\pi^{\prime}(y)\in\textsf{support}(\tau(\pi^{\prime}(x)))$.
It follows that $\pi^{\prime}(y)\smile\textsf{parent}(\pi^{\prime}(x))$ by definition of $F$ and thus $\nu^{\prime}(y)\smile\textsf{parent}(\nu^{\prime}(x))$ also by definition of $F$.
Because $\gamma$ is a morphism,
$\beta_{y}(\sigma^{\prime}(\nu^{\prime}(y),\textsf{parent}(\nu^{\prime}(x)))=\sigma(y,\textsf{parent}(x))$
By definition of $F$ on scenario morphisms:
$\alpha_{y}(\sigma^{\prime}(\pi^{\prime}(y),\textsf{parent}(\pi^{\prime}(x)))=\sigma(y,\textsf{parent}(x))$
By definition of $F$ on games:
$\alpha_{y}\left[\tau(\pi^{\prime}(x))(\pi^{\prime}(y))\right]=\tau(x)(y)$
which proves that the third constraint holds.
Thus, $F(\gamma)$ is a scenario morphism. ∎
Lemma 16.
F preserves the identity morphism.
Proof.
$\text{Id}_{\mathcal{G}}$ is defined as:
  • $\nu^{\prime}=\text{Id}_{\mathcal{N}_{A}}$ (identified with $\text{Id}_{\mathcal{I}_{A}}$ with the abuse of notation)
  • $\beta_{x}=\text{Id}_{\chi(x)}$ for each $x\in\mathcal{N}_{A}$
If we now consider $F(\text{Id}_{\mathcal{G}})$ defined as above and with components $(\pi^{\prime},\alpha)$ then $\pi^{\prime}=\nu^{\prime}=\text{Id}_{\mathcal{I}_{A}}=\text{Id}_{X}$ and for each $x\in X=\mathcal{I}_{A}$, $\alpha_{x}=\beta_{x}=\text{Id}_{\chi(x)}=\text{Id}_{O_{x}}$.
Thus, $F(\text{Id}_{\mathcal{G}})=\text{Id}_{F(\mathcal{G})}$
Lemma 17.
F preserves morphism composition.
Proof.
[…]
Recall that if $\gamma_{1}$ has components $(\nu_{1}^{\prime},\beta_{1})$ and $\gamma_{2}$ has components $(\nu_{2}^{\prime\prime},\beta_{2}^{\prime})$, then the morphism $\gamma_{3}=\gamma_{1}\circ\gamma_{2}:\mathcal{G^{\prime\prime}}\rightarrow\mathcal{G}$ is defined as the tuple $(\nu_{3}^{\prime\prime},\beta_{3})$ where:
  • $\nu_{3}^{\prime\prime}=\nu_{2}^{\prime\prime}\circ\nu_{1}^{\prime}$.
  • $\beta_{3,x}=\beta_{1,x}\circ\beta_{2,\nu_{1}^{\prime}(x)}^{\prime}$ for every $x\in\mathcal{I}_{A}$
$F(\gamma_{1})$ has components $(\pi_{1}^{\prime},\alpha_{1})$ where:
  • $\pi_{1}^{\prime\prime}=\nu_{1}^{\prime}$
  • $\alpha_{1}=\beta_{1}$
$F(\gamma_{2})$ has components ($\pi_{2}^{\prime\prime},\alpha_{2}^{\prime}$) where:
  • $\pi_{2}^{\prime\prime}=\nu_{2}^{\prime\prime}$.
  • $\alpha_{2}^{\prime}=\beta_{2}^{\prime}$.
Thus, by definition of the composition of scenario morphisms, $F(\gamma_{1})\circ F(\gamma_{2})$ has components ($\pi_{3}^{\prime\prime},\alpha_{3}$) where:
  • $\pi_{3}^{\prime\prime}=\pi_{2}^{\prime\prime}\circ\pi_{1}^{\prime}=\nu_{2}^{\prime\prime}\circ\nu_{1}^{\prime}=\nu_{3}^{\prime\prime}$.
  • for any $x\in X=\mathcal{I}_{A}$, $\alpha_{3,x}=\alpha_{1,x}\circ\alpha_{2,\pi_{1}^{\prime}(x)}=\beta_{1,x}\circ\beta_{2,\nu_{1}^{\prime}(x)}=\beta_{3,x}$
This matches exactly the definition of $F(\gamma_{3})$.
Thus $F(\gamma_{3})=F(\gamma_{1}\circ\gamma_{2})=F(\gamma_{1})\circ F(\gamma_{2})$ and F preserves morphism composition.
Theorem 18.
F is a functor.
Proof.
F preserves the identity and composition of morphisms. It is thus a functor from Game to Scenario. ∎
3.3.2 Essential surjectivity
In order to show that $F$ is essentially surjective, we build a functor $G:\textbf{Scenario}\rightarrow\textbf{Game}$. This functor will allow us to show that any scenario is isomorphic to a scenario in the image of Game.